### v2

# Decoherence of Josephson coupling and thermal quenching of the Josephson diode effect in bilayer superconductors

Abstract
Motivated by recent studies on superconducting (SC) diode nonreciprocity, we uncover an unexpected hierarchy of SC-phase decoherence in bilayer superconductors hosting both interlayer Josephson coupling and a Josephson diode effect. Contrary to the conventional single–energy-scale paradigm where Josephson coherence and diode nonreciprocity vanish simultaneously at the SC gap-closing temperature, we demonstrate, using a self-consistent microscopic theory incorporating phase fluctuations, that the system undergoes a sequence of distinct thermal crossovers upon heating: the diode effect disappears first at $T_{\eta}$, Josephson coherence is subsequently lost at $T_{c}$, and the SC gap collapses only at a higher temperature $T_{s}$. Rather than a direct SC–normal transition, the system thus evolves through successive nonreciprocal, reciprocal, and incoherent Josephson regimes before entering the normal state. Counterintuitively, the separation between these regimes is governed not only by interlayer coupling, but also sensitively by in-plane disorder and carrier density. These findings point to a generic hierarchy of SC decoherence in low-dimensional Josephson systems, and suggest broader relevance to layered superconductors, including cuprates and recently discovered nickelates, as well as to SC qubits.
Introduction.—Over the past few decades, Josephson junctions, where two superconductors are coupled through a non-superconducting barrier [1], have emerged not only as the basis of a wide range of superconducting (SC) electronic devices [2, 3, 4], but also as essential building blocks of quantum technologies and quantum computation [5, 6, 7]. In conventional Josephson junctions with a classical or centrosymmetric barrier, SC transport is reciprocal. This reciprocity originates from the current–phase relation (CPR) of the Josephson supercurrent $I_{J}$ as a function of the SC phase difference $\phi$ between the two electrodes [8], which takes the generalized form $I_{J}(\phi)=|J_{1}||\Delta_{1}||\Delta_{2}|\sin(\phi)+2|J_{2}||\Delta_{1}|^{2}|\Delta_{2}|^{2}\sin(2\phi)$. Here, $|J_{1}|$ and $|J_{2}|$ denote the amplitudes of the first and second harmonics, respectively, and $|\Delta|_{1}$ and $|\Delta_{2}|$ are the SC gaps of the two electrodes. Because of this CPR, the critical supercurrent is identical for opposite current directions, reflecting the reciprocal nature of conventional Josephson transport.
Recently, by analogy with ppnn junctions in semiconductor technology which permit low-resistance charge transport in one direction but high resistance in the opposite one, the realization of a Josephson diode effect, i.e., a non-reciprocal critical supercurrent, has attracted considerable attention [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Such an effect enables directional control of dissipationless supercurrents and is of both fundamental interest and practical importance for applications in SC electronics. Realizing the SC diode effect requires the breaking of time-reversal symmetry (and often inversion symmetry) [20], which results in a generalized asymmetric CPR [14, 21, 22],
$I_{J}(\phi)=|J_{1}||\Delta_{1}||\Delta_{2}|\sin(\phi+\alpha)+2|J_{2}||\Delta_{1}|^{2}|\Delta_{2}|^{2}\sin(2\phi),$ (1)
where $\alpha\neq 0$ signals the absence of time-reversal symmetry. As a result, as shown in Fig. 1(b), the maximum supercurrent flowing in the positive direction differs from that in the negative direction, giving rise to non-reciprocal critical currents and the Josephson diode behavior, as experimentally observed in Josephson junction systems with time-reversal symmetry breaking [14]. In particular, with the rapid development of advanced quantum materials, the Josephson diode behavior has been experimentally and theoretically explored in low-dimensional junctions, including a variety of multilayer and bilayer platforms [18, 23, 24, 25, 26, 27, 28, 29, 30, 22]. Such Josephson junctions can host engineered asymmetry, for example arising from twist angles [23, 26, 27, 28, 29, 22, 25], offering a versatile and highly tunable route toward controlling the Josephson diode behavior.
Figure 1: (a) Schematic illustration of a bilayer superconductor with interlayer Josephson coupling. (b) The CPR given by Eq. (1). Inset: Temperature dependence of the critical-current strength (normalized) in the forward and backward directions from Eq. (1) based on the mean-field BCS theory to calculate the SC gap. For simplicity and without loss of generality, we set $\alpha=0.2\pi$ and $|J_{1}|=4|J_{2}||\Delta_{1}||\Delta_{2}|$.
Within the conventional CPR paradigm of Josephson junctions and Josephson diodes [Eq. (1)], both Josephson coherence and diode non-reciprocity, are expected to disappear simultaneously at the characteristic temperature where the SC gap vanishes. In this work, however, we emphasize the central role of SC phase coherence and rigorously demonstrate, within a self-consistent microscopic framework incorporating SC phase fluctuations, that low-dimensional Josephson junctions, such as bilayer superconductors, undergo a sequence of distinct thermal crossovers upon heating (summarized in Fig. 3(a)). Specifically, as temperature increases, the system evolves from a non-reciprocal Josephson phase to a reciprocal phase, followed by an incoherent Josephson state, and finally into the normal state, such that the Josephson diode effect disappears first at a characteristic temperature $T_{\eta}$, followed by the loss of Josephson coherence at $T_{c}$, and finally the vanishing of the SC gap at a higher temperature $T_{s}$. These findings reveal a hierarchy of decoherence processes that are generic to low-dimensional Josephson systems and, more broadly, can potentially be extended to understand layered SC materials, such as cuprates [31, 32, 33, 34, 35, 36, 37, 38] and recently discovered nickelates [39, 40, 41, 42, 43, 44], particularly their $c$-axis transport behaviors.
Model.—We start with a minimal mean-field Hamiltonian description [45, 46] of a bilayer superconductor with interlayer Josephson coupling, as schematically illustrated in Fig. 1(a),
$H=$ $\!\!\sum_{j=1,2}\Big[\!\int\!{d{\bf x}}\,\Psi^{\dagger}_{j}({\bf x})\left(\begin{array}[]{cc}\xi_{j{\hat{\bf p}}}&\Delta_{j}({\bf x})\\ \Delta_{j}^{*}({\bf x})&\xi_{j{\bf{\hat{p}}}}\end{array}\right)\Psi_{j}({\bf x})\!-\!\frac{|\Delta_{j}|^{2}}{U}\Big]$ (4)
$\mbox{}-\frac{1}{2}(J_{1}\Delta_{1}\Delta^{*}_{2}\!+\!c.c.)\!-\!\frac{1}{2}[J_{2}(\Delta_{1}\Delta^{*}_{2})^{2}\!+\!c.c.]\!+\!H_{I},$ (5)
where $\Psi^{\dagger}_{j}=(c^{\dagger}_{j\uparrow},c_{j{\downarrow}})$ is the Nambu spinor in layer $j$ with $c^{\dagger}_{js}$ and $c_{js}$ being the creation and annihilation operators, respectively; $\xi_{j{\bf k}}$ denotes the single-particle dispersion measured from the chemical potential, and $U<0$ is the SC pairing interaction; the momentum operator $\hat{\bf p}=-i\hbar{\nabla}$, while $\bf x$} denotes the spatial coordinate confined to the in-plane space of each layer; $H_{I}$ denotes the long-range Coulomb interactions (Sec. SI). The SC order parameter in each layer reads [45, 46]
$\Delta_{j}({\bf x})=|\Delta_{j}|e^{i\phi_{j}({\bf x})},$ (6)
where $|\Delta_{j}|$ and $\phi_{j}({\bf x})$ are the SC gap and phase, respectively. Here the interlayer Josephson coupling includes both the conventional first harmonic and a higher-order second harmonic contribution, and $J_{1,2}$ are generally complex interlayer Josephson amplitudes [14, 21, 22]. Within the conventional derivation of the Josephson effect [47, 48], these couplings originate from the interlayer hopping of electrons, which mediates the tunneling of Cooper pairs between the two SC layers. While the first harmonic $J_{1}$ arises from the lowest-order tunneling of a single Cooper pair, the second harmonic $J_{2}$ emerges from higher-order tunneling processes, involving correlated tunneling of two Cooper pairs via virtual intermediate states [47].
We consider the symmetric bilayer case and assume identical electronic structures in the two layers, i.e., $\xi_{j{\bf k}}=\xi_{\bf k}$, thereby yielding $|\Delta_{j}|=|\Delta|$. In addition, we consider a complex first-harmonic Josephson coupling $J_{1}=|J_{1}|e^{i\alpha}$, while the second-harmonic coupling is taken to be real, $J_{2}=J_{2}^{*}$. The phase shift $\alpha$ represents a nontrivial relative phase between different Josephson harmonics that explicitly breaks time-reversal symmetry, such that $2{\rm arg}(J_{1})\neq{\rm arg}(J_{2})$, and cannot be removed by a global redefinition of the Josephson phases. Such a phase shift can arise from the interplay of interlayer twist effect [20, 11, 13, 23, 26, 27, 28, 29, 22, 21, 25], which generates multiple inequivalent tunneling channels, and external or internally generated effective magnetic fields that imprint phase frustration onto these channels. The present work further explores the physical implications of the SC phase coherence.
To consider the SC phase coherence, we introduce the relative (Josephson) and total SC phase fields as
$\phi({\bf x})$ $=$ $\phi_{1}({\bf x})-\phi_{2}({\bf x}),$ (7)
$\theta({\bf x})$ $=$ $[\phi_{1}({\bf x})+\phi_{2}({\bf x})]/2.$ (8)
Within path-integral formalism, integrating out the fermionic degrees of freedom leads to an effective action (Sec. SI)
$S_{\rm eff}=S_{\rm intra}\big(|\Delta|,{\bf p}_{s}^{2}\big)+S_{J}(\phi).$ (9)
Here, $S_{\rm intra}$ governs the intralayer SC properties, while ${\bf p}_{s}=\bm{\nabla}\theta({\bf x})/2$ is the physically meaningful, gauge-invariant quantity describing fluctuations of the total phase $\theta({\bf x})$ [49, 50]. These fluctuations correspond to the in-phase collective mode [51, 50, 52], associated with the bosonic Nambu–Goldstone (NG) mode of the SC state [53, 54, 55, 56], and involve long-range Coulomb interactions in the symmetric channel. In 2D systems, this mode remains gapless and an active collective excitation [52, 57, 58, 59, 60], in contrast to the dynamically inert case in 3D [61, 62], where the NG mode is fully gapped out by the Anderson–Higgs mechanism [63]. In our recent works for 2D superconductors [58, 59, 60], we have self-consistently incorporated these finite NG-phase fluctuations into the SC gap equation in the presence of long-range Coulomb interactions. As a result, the SC gap and its disappearance temperature $T_{s}$ acquire nontrivial dependence on disorder and carrier density, reflecting fluctuation-induced renormalization effects beyond mean-field theory [58, 59, 60]. For completeness, we directly adopt this approach for the in-plane sector to determine $|\Delta(T)|$ and $\langle{\bf p}_{s}^{2}\rangle$ (see Sec. SIII), which enter the subsequent interlayer Josephson formulation as external inputs. Nevertheless, the central result derived below follows from general physics and gauge-invariant considerations and does not depend on the specific microscopic realization of these quantities.
The essential effective action for the interlayer Josephson coupling and the out-of-phase collective mode (known as the Leggett-type [64] Josephson phase mode) reads
$S_{J}(\phi)=$ $\!\int\!{dR}\Big\{\frac{D_{l}}{2}\Big(\frac{\partial_{t}\phi}{2}\Big)^{2}-\frac{f_{s}}{4}\Big(\frac{\nabla\phi}{2}\Big)^{2}+|J_{1}||\Delta|^{2}\cos(\phi+\alpha)$
$\mbox{}+|J_{2}||\Delta|^{4}\cos(2\phi)\Big\}.$ (10)
Here, $D_{l}$ denotes the effective density of states renormalized by the interlayer long-range Coulomb interactions, and $f_{s}$ is the phase stiffness, written as (see Sec. SI for detailed derivation)
$f_{s}=f_{s}(|\Delta|^{2},\langle{{\bf p}_{s}^{2}}\rangle)=\!\frac{\mathrm{v}^{2}_{F}|\Delta|^{2}}{1+\xi/l}\sum_{\bf k}\frac{f(E_{\bf k}^{-})-f(E_{\bf k}^{+})}{2E^{3}_{\bf k}},$ (11)
which depends on gap $|\Delta(T)|$ and NG phase fluctuations $\langle{\bf p}_{s}^{2}\rangle$. Here, the fermionic quasiparticle spectrum $E_{\bf k}^{\pm}={\bf v_{\bf k}}\cdot{\bf p}_{s}\pm{E_{\bf k}}$ with $E_{\bf k}=\sqrt{\xi_{\bf k}^{2}+|\Delta|^{2}}$; the prefactor $(1+\xi/l)^{-1}$ accounts for the disorder-induced reduction of the phase stiffness, consistent with Tinkham’s interpolation [65] and subsequent microscopic studies [66, 67, 68, 45], where $\xi=\hbar v_{F}/|\Delta|$ is the SC coherence length and $l=v_{F}\tau$ is the mean free path with $\tau$ being the effective scattering time. Within a parabolic-band approximation, $f_{s}$ reduces to the familiar form ${n_{s}}/m^{*}$ [51, 50, 52, 66, 68], with $n_{s}$ being superfluid density and $m^{*}$ the effective mass.
According to the action $S_{J}(\phi)$, the interlayer Josephson coupling explicitly gaps the out-of-phase mode. We show in the following that the Josephson decoherence proceeds through smooth renormalization. Specifically, we decompose the Josephson phase as $\phi=\phi_{e}+\delta\phi$ where $\delta\phi$ represents Josephson phase fluctuations around equilibrium configuration $\phi_{e}$. With $\langle\delta\phi\rangle=0$, using the cumulant expansion (see Sec. SII. A for detailed derivations)
$\langle\cos(n\phi)\rangle={\rm Re}[e^{in\phi_{e}}\langle{e^{in\delta\phi}}\rangle]=\cos(n\phi_{e})e^{-n^{2}\langle{\delta\phi}^{2}\rangle/2},$ (12)
the thermally averaged Josephson energy reads
$E_{J}(\phi_{e})$ $=$ $-|J_{1}||\Delta|^{2}\big\langle\cos(\phi+\alpha)\big\rangle-|J_{2}||\Delta|^{4}\langle\cos(2\phi)\rangle$ (13)
$=$ $-{\bar{J}}_{1}(T)\cos(\phi_{e}+\alpha)-{\bar{J}}_{2}(T)\cos(2\phi_{e}),$
where the effective Josephson couplings are renormalized by Josephson phase fluctuations as
${\bar{J}}_{1}(T)$ $=$ $|J_{1}||\Delta(T)|^{2}\exp[-\langle\delta\phi^{2}(T)\rangle/2],$ (14)
${\bar{J}}_{2}(T)$ $=$ $|J_{2}||\Delta(T)|^{4}\exp[-2\langle\delta\phi^{2}(T)\rangle].$ (15)
The optimal equilibrium phase $\phi^{\rm op}_{e}$ is determined by global minimizer of the fluctuation-renormalized Josephson energy
$\phi_{e}^{\rm op}=\arg\min_{\phi^{e}\in[0,2\pi)}E_{J}(\phi_{e}).$ (16)
The Josephson phase fluctuations can be derived within the quantum statistic mechanism based on the equation of motion of Josephson phase fluctuations $\delta\phi$ obtained from the action $S_{J}(\phi)$ (Sec. SII. B), and their average is derived as (Sec. SII. C)
$\langle\delta\phi^{2}(T)\rangle=\int\frac{2d{\bf q}}{(2\pi)^{2}}\frac{2n_{B}[\omega_{\rm L}({q})]+1}{D_{l}\omega_{\rm L}({q})},$ (17)
which corresponds to the standard bosonic excitation, consisting of contributions from both thermal excitations $2n_{B}(\omega_{\rm L})$ (thermal fluctuations) and zero-point oscillations (quantum fluctuations). Here, the function $n_{B}(x)$ is the Bose distribution. The excitation spectrum $\omega_{\rm L}(q)$ of this Leggett-type [64] collective phase mode is determined by
$D_{l}\omega^{2}_{\rm L}({q})={4{\bar{J}}_{1}\cos(\phi^{\rm op}_{e}\!+\!\alpha)}+{16{\bar{J}}_{2}\cos(2\phi^{\rm op}_{e})}+{f_{s}q^{2}}/{2},$ (18)
corresponding to a massive (gapped) collective excitation.
Consequently, Eqs. (13)–(18) form a nontrivial closed set of self-consistent equations, which determine the effective Josephson couplings ${\bar{J}}_{1}(T)$ and ${\bar{J}}_{2}(T)$ renormalized by the Josephson phase fluctuations. Further considering the Josephson supercurrent injection $I_{J}$ (Sec. SII. A), this procedure yields a qualitatively modified Josephson CPR,
$I_{J}(\phi_{e})=2e[{\bar{J}}_{1}(T)\sin(\phi_{e}+\alpha)+2{\bar{J}}_{2}(T)\sin(2\phi_{e})].$ (19)
Consequently, the Debye–Waller–like factors [69, 70] appearing in the Josephson couplings, $\exp({-\langle{\delta\phi^{2}}(T)\rangle}/2)$ in ${\bar{J}}_{1}(T)$ and $\exp({-2\langle{\delta\phi^{2}}(T)\rangle})$ in ${\bar{J}}_{2}(T)$, which originate from Josephson phase fluctuations, render the temperature dependence of the Josephson CPR no longer governed solely by a single energy scale set by the SC gap as in Eq. (1). These factors, in fact, encode the decoherence of the Josephson coupling induced by SC phase fluctuations, yielding a Josephson response that is governed by both the SC gap and Josephson phase coherence.
In particularly, in the limit where $\bar{J}_{1}\rightarrow 0$ and $\bar{J}_{2}\rightarrow 0$ as a consequence of signigicant Josephson phase fluctuations, while the SC gap $|\Delta|$ remains finite, the equilibrium Josephson phase can no longer be uniquely determined from the minimum of the Josephson energy $E_{J}$ in Eq. (13). This indicates that SC phase coherence across the Josephson junction is not established, despite the presence of a finite pairing gap.
Figure 2: Temperature dependence of the Josephson critical current, the diode efficiency and the SC gap self-consistently calculated by (a) mean-field and (b) phase-fluctuation theories. We set $\tau|\Delta_{\rm MF}(0)|=0.2$, $v_{F}=10^{5}$~m/s and $|\Delta_{\rm MF}(0)|=2$~meV, where $|\Delta_{\rm MF}(0)|$ is the zero-temperature gap from mean-field theory.
Results.—We perform fully self-consistent numerical calculations, incorporating both the intralayer SC properties $|\Delta(T)|$ and $\langle{p_{s}^{2}}(T)\rangle$, and the proposed Josephson-related framework in this study [Eqs. (13)–(18)]. The specific model parameters used in simulation are provided in the Supplemental Material (Sec. SIV). Following previous works [14, 21, 22], we define the Josephson critical current as $I_{c}(T)={\rm max}(|I_{c}^{+}(T)|,|I_{c}^{-}(T)|)$, and introduce the diode efficiency
$\eta(T)=\Bigg|\frac{|I_{c}^{+}(T)|-|I_{c}^{-}(T)|}{|I_{c}^{+}(T=0)|+|I_{c}^{-}(T=0)|}\Bigg|$ (20)
which quantifies the relative strength of the nonreciprocal Josephson response at finite temperatures, normalized to its $T=0$ value. Here, $I_{c}^{+}(T)$ and $I_{c}^{-}(T)$ denote the critical Josephson currents for positive and negative current directions, respectively. For comparison, results obtained with and without phase fluctuations are shown in Fig. 2. It is noted that even at $T=0$, including the contribution from the zero-point Josephson phase fluctuations suppresses the zero-temperature diode efficiency $\eta(0)$, while the NG and Josephson phase fluctuations also reduce the zero-temperature gap $|\Delta(0)|$ (as discussed in Refs. [58, 59, 60]) and the critical current $I_{c}(0)$ (see Sec. SV for more details related to zero-point oscillations), respectively. This zero-point effect suggests the intrinsic quantum sensitivity of low-dimensional Josephson-diode systems, and it represents a fundamental quantum limitation that cannot be circumvented by cooling. Quantum fluctuations therefore impose an intrinsic limit on Josephson coherence and nonreciprocal transport in low-dimensional Josephson systems.
For the essential critical behaviors, in the absence of phase fluctuations, as shown in Fig. 2(a), both SC and Josephson properties are solely determined by the SC gap, such that the Josephson critical current $I_{c}(T)$ and the diode efficiency $\eta(T)$ vanish at the same temperature at which the SC gap $|\Delta(T)|$ closes. In contrast, when significant Josephson phase fluctuations are activated, as shown in Fig. 2(b), a clear separation of energy scales emerges upon increasing temperature: the diode efficiency $\eta(T)$ disappears first, followed by the suppression of the Josephson critical current $I_{c}(T)$, while the SC gap $|\Delta(T)|$ remains finite and vanishes only at a higher temperature.
This is because that when Josephson phase fluctuations are activated, the Debye–Waller–like factors in ${\bar{J}}_{1}$ and ${\bar{J}}_{2}$ lead to a rapid suppression of the effective Josephson couplings as the fluctuation strength increases with temperature. As a result, the Josephson coupling is exponentially reduced by Josephson phase fluctuations, even in a regime where the SC gap remains finite, causing the Josephson effect to vanish at a temperature $T_{c}$ lower than the gap-closing temperature $T_{s}$. Importantly, higher-order Josephson couplings are suppressed more strongly than the lowest-order term: the second-harmonic coupling ${\bar{J}}_{2}\propto\exp({-2\langle{\delta\phi^{2}}\rangle})$ decays parametrically faster than the first-harmonic coupling ${\bar{J}}_{1}\propto\exp({-\langle{\delta\phi^{2}}\rangle}/2)$ as phase fluctuations increase with temperature. This hierarchy implies that non-sinusoidal components of the Josephson current–phase relation are particularly fragile against phase decoherence, with higher-order harmonics (e.g., $n\geq 3$) expected to be even more fragile according to Eq. (12). Since the Josephson diode nonreciprocity heavily relies on these non-sinusoidal components [14, 21, 22], the Josephson diode effect is expected to disappear at a temperature $T_{\eta}$ lower than the critical temperature $T_{c}$ at which the Josephson critical current is suppressed.
Figure 3: (a) Schematic illustration of phase diagram upon heating. (b) Temperature dependence of the critical-current strength in the forward and backward directions. The distinct temperatures $T_{\eta}$, $T_{c}$, and $T_{s}$ under reducing phase stiffness by (c) increasing disorder at $v_{F}=v_{0}$ and (d) reducing the Fermi velocity at $\tau|\Delta_{\rm MF}(0)|=0.1$. Here, $v_{0}=10^{5}$~m/s and $|\Delta_{\rm MF}(0)|=2$~meV, where $|\Delta_{\rm MF}(0)|$ is the zero-temperature gap calculated from mean-field theory.
The emergence of three distinct temperature scales suggests the crucial role of Josephson-phase coherence in stabilizing nonreciprocal Josephson transport in low-dimensional SC systems, allowing Josephson phenomena to disappear while the SC gap remains finite. Accordingly, as schematically summarized in Fig. 3(a), the system undergoes a sequence of distinct thermal crossovers upon heating: it first evolves from a nonreciprocal Josephson phase to a reciprocal Josephson phase at a characteristic temperature $T_{\eta}$, followed by an incoherent Josephson state at $T_{c}$, above which phase coherence across the Josephson junction is lost while the SC gap remains finite, and finally enters the normal state at $T_{s}$.
We emphasize that the renormalized Josephson couplings in Eqs. (14) and (15) are derived in a mathematically rigorous manner, without reliance on model-specific parameters or physical assumptions, and are therefore generic. Then, the key point is whether Josephson phase fluctuations can be effectively excited. Typically, the Josephson coupling is weak, such that the gap of the Josephson phase mode $\omega_{L}(q=0)$ in Eq. (18)] lies below the SC gap, rendering it a thermally active low-energy excitation. Notably, an equally important factor is the phase stiffness $f_{s}$, which determines the available phase space for bosonic phase fluctuations entering the fluctuation integrals [Eq. (17)]. In bulk systems, the phase stiffness is generally large, strongly suppressing phase fluctuations. By contrast, in low-dimensional systems the phase stiffness is significantly reduced, owing both to a smaller Fermi velocity and to the intrinsically enhanced disorder associated with reduced dimensionality [71, 72]. Increasing disorder and reducing the Fermi velocity, as shown in Figs. 3(c) and (d), respectively, lead to a suppression of $T_{s}$, $T_{c}$, $T_{\eta}$, while markedly increasing the separation among these characteristic temperatures. As a result, the hierarchy of SC decoherence scales becomes increasingly pronounced. Among these effects, the suppression of gap closing temperature $T_{s}$ originates from the enhanced NG phase fluctuations (as discussed in Refs. [58, 59, 60]), whereas all the remaining Josephson features are primarily driven by the enhanced Josephson phase fluctuations (i.e., decoherence).
Discussion.—Beyond the conventional single-energy-scale paradigm of the Josephson CPR, the present study demonstrates that phase decoherence constitutes an independent and more fragile energy scale in low-dimensional Josephson systems, and hence, Josephson physics itself in this case splits into multiple decoherence regimes, including diode nonreciprocity as the most fragile scale. We therefore predict that the forward and backward critical currents in low-dimensional Josephson-diode systems, $|I_{c}^{+}(T)|$ and $|I_{c}^{-}(T)|$, to merge as temperature increases, before vanishing altogether at $T_{c}$, as shown in Fig. 3(b). A similar tendency appears to be present in a wide range of recent experiments (e.g., Refs. [18, 19]), where the data exhibit a temperature evolution qualitatively consistent with that shown in Fig. 3(b).
It should be emphasized that the vanishing of the Josephson critical current and the disappearance of the diode effect correspond to exponential thermal crossovers, rather than conventional thermodynamic phase transitions of first or second order. Both the diode efficiency $\eta(T)$ and the Josephson critical current $I_{c}(T)$ shown in Fig 2(b) exhibit behaviors with continuous temperature derivatives in the vicinity of their respective disappearance temperatures, in contrast to the mean-field scenario shown in Fig 2(a), where discontinuities in temperature derivatives are expected. Consequently, these crossovers do not exhibit thermodynamic signatures such as latent heat or a jump in the specific heat, in contrast to the SC gap-closing transition, which is a genuine second-order phase transition [45, 46]. Such critical behaviors of the Josephson properties therefore constitute a key distinguishing feature between mean-field and fluctuation-based descriptions of Josephson systems.
The present work provides a tractable description of Josephson decoherence. Owing to the structural similarity [51, 60], this tractable framework can be naturally extended to layered superconductors with multilayer structures and Josephson coupling $J_{c}\cos(\theta_{n+1}-\theta_{n})$ between different layers (where $\theta_{n}$ is the SC phase in the $n$-th layer), such as cuprates [31, 32, 33, 34, 35, 36, 37, 38] and recently discovered nickelates [39, 40, 41, 42, 43, 44], particularly for their $c$-axis transport behavior. In this context, one can anticipate the disappearance of $c$-axis superconductivity (zero-resistance) while the SC gap persists above a characteristic critical temperature. Moreover, owing to the close analogy in the underlying derivation of the excitations, similar considerations also apply to SC transmon qubits [73, 74, 75]. Our results indicate that higher-order Josephson harmonics are increasingly fragile against phase decoherence. This effect is expected to be relevant for the stability/coherence of Josephson-based SC qubits, where such higher harmonics play an essential role in determining the qubit spectrum and anharmonicity.
Acknowledgments.—This work is supported by the US Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of Computational Materials Sciences Program. F.Y. and L.Q.C. also appreciate the generous support from the Donald W. Hamer Foundation through a Hamer Professorship at Penn State.