### v3

# Superconducting Decoherence and Thermal Quenching of the Josephson Diode Effect in Low-Dimensional Josephson Systems

Abstract
Motivated by recent studies on superconducting (SC) diode nonreciprocity, we uncover a generic smooth SC-phase decoherence mechanism in low-dimensional Josephson structures. Contrary to the conventional single-energy-scale paradigm where Josephson coherence and diode nonreciprocity vanish simultaneously only at the SC gap-closing temperature, we demonstrate, within a fully self-consistent microscopic framework beyond mean-field theory, that SC phase fluctuations generically split these phenomena into distinct energy scales. As a result, rather than a single SC-normal transition, the system exhibits a sequence of distinct thermal crossovers upon heating: the diode effect disappears first at $T_{\eta}$, Josephson coherence is subsequently lost at $T_{c}$, and the SC gap collapses only at a higher temperature $T_{s}$. Using a bilayer SC system as a concrete example, we show that the separation between these temperature scales is not solely dictated by Josephson coupling, but is instead strongly and counterintuitively shaped by the in-plane disorder and carrier density. These findings reveal that smooth SC phase decoherence introduces a distinct and more fragile energy scale, with potential implications for layered superconductors such as cuprates and recently discovered nickelates, as well as for SC qubit platforms.
Introduction.—Over the past few decades, Josephson junctions, where two superconductors are coupled through a non-superconducting barrier [1], have emerged not only as the basis of a wide range of superconducting (SC) electronic devices [2, 3, 4], but also as essential building blocks of quantum technologies and quantum computation [5, 6, 7]. The physics of Josephson systems is governed by the current-phase relation (CPR) of the Josephson supercurrent $I_{J}$ as a function of the SC phase difference $\phi$ between the two electrodes [8], which acquires a generalized form [9, 10, 11] upon a standard perturbative treatment [12, 13] of the electron tunneling amplitude:
$I_{J}(\phi)=|J_{1}||\Delta_{1}||\Delta_{2}|\sin(\phi+\alpha)+2|J_{2}||\Delta_{1}|^{2}|\Delta_{2}|^{2}\sin(2\phi).$ (1)
Here $|J_{1}|$ and $|J_{2}|$ denote the amplitudes of the first and second harmonics, and $|\Delta_{1,2}|$ are the SC gaps of the two electrodes. The phase shift $\alpha$ is a gauge-invariant quantity that cannot be removed by a gauge transformation. In conventional Josephson junctions with a classical or centrosymmetric barrier, $\alpha=0$, and the CPR is symmetric, leading to identical critical supercurrents for opposite current directions and therefore reciprocal SC transport. In contrast, $\alpha\neq 0$ signals the absence of time-reversal symmetry [14], typically induced by internal or external magnetism, resulting in an asymmetric CPR [as shown in Fig. 1(b)]. This asymmetry gives rise to the Josephson diode effect, namely a nonreciprocal critical supercurrent.
By analogy with semiconductor ppnn junctions that allow low-resistance charge transport in one direction but high resistance in the opposite direction, Josephson diodes enable directional SC transport, and has therefore attracted considerable attention recently [15, 16, 17, 18, 19, 9, 20, 21, 22, 23]. Owing to the rapid development of advanced quantum materials, the Josephson diode behavior has been experimentally and theoretically explored in low-dimensional junctions, including a variety of multilayer and bilayer platforms [24, 25, 26, 27, 28, 29, 30, 31, 32, 11]. Such Josephson junctions can host engineered asymmetry, for example arising from twist angles [25, 28, 29, 30, 31, 11, 27], offering a versatile and highly tunable route toward controlling the Josephson diode behavior.
Figure 1: (a) Schematic illustration of a low-dimensional Josephson system. (b) The CPR given by Eq. (1). Inset: Critical-current strength (normalized) in the forward and backward directions from Eq. (1) based on the mean-field BCS theory to calculate SC gap. Without loss of generality, we set $\alpha=0.2\pi$, $|\Delta_{1}|=|\Delta_{2}|$, and $|J_{1}|=4|J_{2}||\Delta_{1}||\Delta_{2}|$.
Within the conventional paradigm, all Josephson phenomena, including phase coherence and diode nonreciprocity, are expected to disappear only when the SC gap collapses, even in low dimensions, since the Berezinskii-Kosterlitz-Thouless (BKT) mechanism, the standard mechanism for destroying SC phase coherence through vortex unbinding, does not apply to the Josephson phase that is locked by Josephson couplings. Here we demonstrate, within a self-consistent microscopic framework beyond mean-field theory, that smooth SC phase decoherence generically introduces a distinct and more fragile energy scale in low-dimensional Josephson systems such as bilayer superconductors. Consequently, Josephson phase coherence and hence all Josephson effects disappears well before the SC gap collapses. Strikingly, the Josephson diode nonreciprocity vanishes at an even lower temperature, preceding the loss of Josephson coherence itself. This reveals a generic SC-decoherence mechanism in low-dimensional Josephson systems, distinct from the BKT mechanism, and instead of a single SC-normal transition, the system undergoes a sequence of thermal crossovers upon heating prior to the SC-gap collapse.
Model.—We start with an effective Hamiltonian description [33, 34], taking a bilayer superconductor with interlayer Josephson coupling as an example [as illustrated in Fig. 1(a)],
$H=$ $\!\!\sum_{j=1,2}\Big[\!\int\!{d{\bf x}}\,\Psi^{\dagger}_{j}({\bf x})\left(\begin{array}[]{cc}\xi_{j{\hat{\bf p}}}&\Delta_{j}({\bf x})\\ \Delta_{j}^{*}({\bf x})&\xi_{j{\bf{\hat{p}}}}\end{array}\right)\Psi_{j}({\bf x})\!-\!\frac{|\Delta_{j}|^{2}}{U}\Big]$ (4)
$\mbox{}-\frac{1}{2}(J_{1}\Delta_{1}\Delta^{*}_{2}\!+\!c.c.)\!-\!\frac{1}{2}[J_{2}(\Delta_{1}\Delta^{*}_{2})^{2}\!+\!c.c.]\!+\!H_{I},$ (5)
where $\Psi^{\dagger}_{j}=(c^{\dagger}_{j\uparrow},c_{j{\downarrow}})$ is the Nambu spinor in layer $j$ with $c^{\dagger}_{js}$ and $c_{js}$ being the creation and annihilation operators, respectively; $\xi_{j{\bf k}}$ denotes the single-particle dispersion measured from the chemical potential, and $U<0$ is the SC pairing interaction; the momentum operator $\hat{\bf p}=-i\hbar{\nabla}$, while $\bf x$} denotes the spatial coordinate confined to the in-plane space of each layer; $H_{I}$ denotes the long-range Coulomb interactions (see Sec. SI). The SC order parameter in each layer reads [33, 34]
$\Delta_{j}({\bf x})=|\Delta_{j}|e^{i\phi_{j}({\bf x})},$ (6)
where $|\Delta_{j}|$ and $\phi_{j}({\bf x})$ are the SC gap and phase, respectively. We consider the symmetric bilayer case and assume identical electronic structures in the two layers, i.e., $\xi_{j{\bf k}}=\xi_{\bf k}$, thereby yielding $|\Delta_{j}|=|\Delta|$. The interlayer Josephson coupling here includes both the conventional first harmonic and a higher-order second harmonic contribution [9, 10, 11]. Within the conventional derivation of the Josephson effect  [12, 13], these couplings originate from the interlayer hopping of electrons, which mediates the Cooper-pairs tunneling between the two SC layers via virtual intermediate states. The first harmonic $J_{1}$ arises from the lowest-order tunneling of a single Cooper pair, and the second harmonic $J_{2}$ emerges from correlated tunneling processes of two Cooper pairs. These Josephson amplitudes are generally complex. By gauge transformation, one can take a complex $J_{1}=|J_{1}|e^{i\alpha}$, while taking $J_{2}$ to be real, $J_{2}=J_{2}^{*}$. The phase shift $\alpha$ represents a nontrivial relative phase between different Josephson harmonics that explicitly breaks time-reversal symmetry, such that $2{\rm arg}(J_{1})\neq{\rm arg}(J_{2})$, and cannot be removed by a global redefinition of the Josephson phases. Such a phase shift can arise from the interplay of interlayer twist effect [14, 17, 19, 25, 28, 29, 30, 31, 11, 10, 27], which generates multiple inequivalent tunneling channels, and external or internally generated effective magnetic fields that imprint phase frustration onto these channels.
To consider the SC phase coherence, we introduce the relative (Josephson) and total SC phase fields as
$\phi({\bf x})$ $=$ $\phi_{1}({\bf x})-\phi_{2}({\bf x}),$ (7)
$\theta({\bf x})$ $=$ $[\phi_{1}({\bf x})+\phi_{2}({\bf x})]/2.$ (8)
Within path-integral formalism, integrating out the fermionic degrees of freedom leads to an effective action (Sec. SI)
$S_{\rm eff}=S_{\rm intra}\big(|\Delta|,{\bf p}_{s}^{2}\big)+S_{J}(\phi).$ (9)
Here, $S_{\rm intra}$ governs the intralayer SC properties, while ${\bf p}_{s}=\bm{\nabla}\theta({\bf x})/2$ is the physically meaningful, gauge-invariant quantity describing fluctuations of the total phase $\theta({\bf x})$ [35, 36], which correspond to the in-phase collective mode [37, 36, 38], associated with the bosonic Nambu–Goldstone (NG) mode of the SC state [39, 40, 41, 42], and involve the long-range Coulomb interactions in the symmetric channel (see Sec. SI).
The microscopically derived effective action for the interlayer Josephson coupling and the out-of-phase collective mode (known as the Leggett-type [43] Josephson phase mode) reads
$S_{J}(\phi)=$ $\!\int\!{dR}\Big[\frac{D_{l}}{2}\Big(\frac{\partial_{t}\phi}{2}\Big)^{2}\!-\!\frac{f_{s}}{4}\Big(\frac{\nabla\phi}{2}\Big)^{2}\!+\!|J_{1}||\Delta|^{2}\cos(\phi\!+\!\alpha)$
$\mbox{}+|J_{2}||\Delta|^{4}\cos(2\phi)\Big].$ (10)
Here, $D_{l}$ denotes the effective density of states renormalized by the long-range Coulomb interactions in the anti-symmetric channel (see Sec. SI), and $f_{s}$ is the phase stiffness. Clearly, the interlayer Josephson coupling explicitly gaps the out-of-phase mode. We decompose the Josephson phase as $\phi=\phi_{e}+\delta\phi$ where $\delta\phi$ represents Josephson phase fluctuations around equilibrium configuration $\phi_{e}$. With $\langle\delta\phi\rangle=0$, using the cumulant expansion (see Sec. SII. A for detailed derivations)
$\langle\cos(m\phi)\rangle\!=\!{\rm Re}[e^{im\phi_{e}}\langle{e^{im\delta\phi}}\rangle]\!=\!{\rm Re}\Big[e^{im\phi_{e}}\!\!\sum_{n=0}\frac{\langle(im\delta\phi)^{2n}\rangle}{(2n)!}\Big]$
$={\rm Re}\Big[e^{im\phi_{e}}\sum_{n=0}\frac{(2n-1)!!(-1)^{n}m^{2n}\langle\delta\phi^{2}\rangle^{n}}{(2n)!}\Big]$
$={\rm Re}[e^{im\phi_{e}}e^{-\langle{\delta\phi}^{2}\rangle/2}]=\cos(m\phi_{e})e^{-m^{2}\langle{\delta\phi}^{2}\rangle/2},$ (11)
the thermally averaged Josephson energy reads
$E_{J}(\phi_{e})$ $=$ $-|J_{1}||\Delta|^{2}\big\langle\cos(\phi+\alpha)\big\rangle-|J_{2}||\Delta|^{4}\langle\cos(2\phi)\rangle$ (12)
$=$ $-{\bar{J}}_{1}(T)\cos(\phi_{e}+\alpha)-{\bar{J}}_{2}(T)\cos(2\phi_{e}),$
where the effective Josephson couplings are renormalized by Josephson phase fluctuations as
${\bar{J}}_{1}(T)$ $=$ $|J_{1}||\Delta(T)|^{2}\exp[-\langle\delta\phi^{2}(T)\rangle/2],$ (13)
${\bar{J}}_{2}(T)$ $=$ $|J_{2}||\Delta(T)|^{4}\exp[-2\langle\delta\phi^{2}(T)\rangle].$ (14)
The optimal equilibrium phase $\phi^{\rm op}_{e}$ is determined by global minimizer of the fluctuation-renormalized Josephson energy
$\phi_{e}^{\rm op}=\arg\min_{\phi^{e}\in[0,2\pi)}E_{J}(\phi_{e}).$ (15)
The Josephson phase fluctuations can be derived within the quantum statistic mechanism based on the equation of motion of Josephson phase fluctuations $\delta\phi$ obtained from the action $S_{J}(\phi)$ (Sec. SII. B), and their average is derived as (Sec. SII. C)
$\langle\delta\phi^{2}(T)\rangle=\int\frac{2d{\bf q}}{(2\pi)^{2}}\frac{2n_{B}[\omega_{\rm L}({q})]+1}{D_{l}\omega_{\rm L}({q})},$ (16)
which corresponds to the standard bosonic excitation, consisting of contributions from both thermal excitations $2n_{B}(\omega_{\rm L})$ (thermal fluctuations) and zero-point oscillations (quantum fluctuations). Here, the function $n_{B}(x)$ is the Bose distribution. The excitation spectrum $\omega_{\rm L}(q)$ of this Leggett-type [43] collective phase mode is determined by
$D_{l}\omega^{2}_{\rm L}({q})={4{\bar{J}}_{1}\cos(\phi^{\rm op}_{e}\!+\!\alpha)}+{16{\bar{J}}_{2}\cos(2\phi^{\rm op}_{e})}+{f_{s}q^{2}}/{2},$ (17)
corresponding to a massive (gapped) collective excitation.
Consequently, Eqs. (12)–(17) form a nontrivial closed set of self-consistent equations, which determine the effective Josephson couplings ${\bar{J}}_{1}(T)$ and ${\bar{J}}_{2}(T)$ renormalized by the Josephson phase fluctuations. Further considering the Josephson supercurrent injection $I_{J}$ (Sec. SII. A), this procedure yields a qualitatively modified Josephson CPR,
$I_{J}(\phi_{e})=2e[{\bar{J}}_{1}(T)\sin(\phi_{e}+\alpha)+2{\bar{J}}_{2}(T)\sin(2\phi_{e})].$ (18)
The Debye-Waller-like factors appearing in the Josephson couplings, $\exp({-\langle{\delta\phi^{2}}(T)\rangle}/2)$ in ${\bar{J}}_{1}(T)$ and $\exp({-2\langle{\delta\phi^{2}}(T)\rangle})$ in ${\bar{J}}_{2}(T)$, which originate from Josephson phase fluctuations, render the Josephson CPR no longer governed solely by a single energy scale set by the SC gap. These factors encode the decoherence of the Josephson coupling, and the resulting Josephson response is governed jointly by the SC gap and phase coherence. It should be emphasized that this decoherence mechanism is fundamentally different from BKT-type transitions: it does not involve vortex unbinding, topological defects of the global phase, or the characteristic universal jump associated with the BKT transition. Instead, it arises from the smooth decoherence of relative-phase locking between SC components.
In particularly, in the limit where $\bar{J}_{1}\rightarrow 0$ and $\bar{J}_{2}\rightarrow 0$ as a consequence of signigicant Josephson phase fluctuations, while the SC gap $|\Delta|$ remains finite, the equilibrium Josephson phase can no longer be uniquely determined from the minimum of the Josephson energy $E_{J}$ in Eq. (12). This indicates that SC phase coherence across the Josephson junction is not established, despite the presence of a finite SC pairing gap.
Results.—We perform fully self-consistent numerical calculations of the proposed framework [Eqs. (12)–(17)]. The phase stiffness $f_{s}$ is written as (see Sec. SI for detailed derivation)
$f_{s}=\!\frac{\mathrm{v}^{2}_{F}|\Delta|^{2}}{1+\xi/l}\sum_{\bf k}\frac{f(E_{\bf k}^{-})-f(E_{\bf k}^{+})}{2E^{3}_{\bf k}},$ (19)
which depends on gap $|\Delta(T)|$ and NG phase fluctuations $\langle{\bf p}_{s}^{2}\rangle$. Here, the fermionic quasiparticle spectrum $E_{\bf k}^{\pm}={\bf v_{\bf k}}\cdot{\bf p}_{s}\pm{E_{\bf k}}$ with $E_{\bf k}=\sqrt{\xi_{\bf k}^{2}+|\Delta|^{2}}$; the prefactor $(1+\xi/l)^{-1}$ accounts for the disorder-induced reduction of the phase stiffness, where $\xi=\hbar v_{F}/|\Delta|$ is the SC coherence length and $l=v_{F}\tau$ is the mean free path with $\tau$ being the effective scattering time. Within a parabolic-band approximation, $f_{s}$ reduces to the familiar form ${n_{s}}/m^{*}$ [37, 36, 38, 44, 45], with $n_{s}$ being superfluid density and $m^{*}$ the effective mass. In 2D systems, the NG mode remains gapless and an active excitation [38, 46, 47, 48, 49], in contrast to the dynamically inert case in 3D [50, 51], where the NG mode is fully gapped out by the Anderson–Higgs mechanism [52]. Owing to this active nature, the SC gap and its disappearance temperature $T_{s}$ acquire nontrivial dependence on disorder and carrier density, reflecting fluctuation-induced renormalization effects beyond mean-field theory [47, 48, 49]. For completeness, we adopt the developed approach in Refs. [47, 48, 49] for the in-plane sector to determine $|\Delta(T)|$ and $\langle{\bf p}_{s}^{2}\rangle$ (see Sec. SIII), which enter the phase stiffness for interlayer Josephson formulation. The specific model parameters used in simulation are provided in the Supplemental Material (Sec. SIV).
Following previous works [9, 10, 11], we define the Josephson critical current as $I_{c}(T)={\rm max}(|I_{c}^{+}(T)|,|I_{c}^{-}(T)|)$, and introduce the diode efficiency:
$\eta(T)=\Bigg|\frac{|I_{c}^{+}(T)|-|I_{c}^{-}(T)|}{|I_{c}^{+}(T=0)|+|I_{c}^{-}(T=0)|}\Bigg|$ (20)
which quantifies the relative strength of the nonreciprocal Josephson response at finite temperatures, normalized to its $T=0$ value. Here, $I_{c}^{+}(T)$ and $I_{c}^{-}(T)$ denote the critical Josephson currents for positive and negative current directions, respectively. For comparison, results obtained with and without phase fluctuations are shown in Fig. 2. It is noted that even at $T=0$, including the contribution from the zero-point Josephson phase fluctuations suppresses the zero-temperature diode efficiency $\eta(0)$, while the NG and Josephson phase fluctuations also reduce the zero-temperature gap $|\Delta(0)|$ (as discussed in Refs. [47, 48, 49]) and the critical current $I_{c}(0)$ (see Sec. SV for more details related to zero-point oscillations), respectively. This zero-point effect suggests the intrinsic quantum sensitivity of low-dimensional Josephson-diode systems, and it represents a fundamental quantum limitation that cannot be circumvented by cooling. Quantum fluctuations therefore impose an intrinsic limit on Josephson coherence and nonreciprocal transport in low-dimensional Josephson systems.
Figure 2: Temperature dependence of the Josephson critical current, the diode efficiency and the SC gap calculated by (a) mean-field and (b) phase-fluctuation theories. We set $v_{F}=10^{5}$~m/s, l=3.3l=3.3~nm, and a BCS constant of $0.55$. The phase-fluctuation results in (b) reveal a sequence of distinct regimes upon heating: rather than a single SC-normal transition predicted by the mean-field theory in (a), the system upon heating first evolves from a nonreciprocal (NR) Josephson phase to a reciprocal (R) Josephson phase at $T_{\eta}$, followed by a phase-incoherent (PI) Josephson state at $T_{c}$, above which phase coherence across the Josephson junction is lost while the SC gap remains finite, and finally enters the normal (N) state at $T_{s}$.
For the essential critical behaviors, in the absence of phase fluctuations, as shown in Fig. 2(a), both SC and Josephson properties are solely determined by the SC gap, such that the Josephson critical current $I_{c}(T)$ and the diode efficiency $\eta(T)$ vanish at the same temperature at which the SC gap $|\Delta(T)|$ closes. In contrast, when significant Josephson phase fluctuations are activated, as shown in Fig. 2(b), a clear separation of energy scales emerges upon increasing temperature: the diode efficiency $\eta(T)$ disappears first, followed by the suppression of the Josephson critical current $I_{c}(T)$, while the SC gap $|\Delta(T)|$ remains finite and vanishes only at a higher temperature.
This is because that when Josephson phase fluctuations are activated, the Debye–Waller–like factors in ${\bar{J}}_{1}$ and ${\bar{J}}_{2}$ lead to a rapid suppression of the effective Josephson couplings as the fluctuation strength increases with temperature. As a result, the Josephson coupling is exponentially reduced by Josephson phase fluctuations, even in a regime where the SC gap remains finite, causing the Josephson effect to vanish at a temperature $T_{c}$ lower than the gap-closing temperature $T_{s}$. Importantly, higher-order Josephson couplings are suppressed more strongly than the lowest-order term: the second-harmonic coupling ${\bar{J}}_{2}\propto\exp({-2\langle{\delta\phi^{2}}\rangle})$ decays parametrically faster than the first-harmonic coupling ${\bar{J}}_{1}\propto\exp({-\langle{\delta\phi^{2}}\rangle}/2)$ as phase fluctuations increase with temperature. This hierarchy implies that non-sinusoidal components of the Josephson current–phase relation are particularly fragile against phase decoherence, with higher-order harmonics (e.g., $n\geq 3$) expected to be even more fragile according to Eq. (Superconducting Decoherence and Thermal Quenching of the Josephson Diode Effect in Low-Dimensional Josephson Systems). Since the Josephson diode nonreciprocity heavily relies on these non-sinusoidal components [9, 10, 11], the Josephson diode effect is expected to disappear at a temperature $T_{\eta}$ lower than the critical $T_{c}$ at which the Josephson critical current is suppressed.
Figure 3: Distinct temperatures $T_{\eta}$, $T_{c}$, and $T_{s}$ as the phase stiffness is reduced by (a) increasing disorder at $v_{F}=v_{0}$ and (b) decreasing the Fermi velocity at $\tau=0.1\tau_{0}$. Here $v_{0}=10^{5}~\mathrm{m/s}$ and $\tau_{0}v_{0}=33$~nm. For comparison, we also show the mean-field BCS prediction (dashed curves), where $T_{s}^{\rm MF}=T_{\eta}^{\rm MF}=T_{c}^{\rm MF}$, which reduces to a constant independent of disorder and carrier density in 2D. (c) The theoretical critical currents $|I_{c}^{+}(T)|$ and $|I_{c}^{-}(T)|$ for $\tau=0.2\tau_{0}$ and $v_{F}=v_{0}$. For comparison, we present related experimental data in (d) and (e), extracted from Refs. [23] and [24], respectively. The qualitative merging of the two critical currents with increasing $T$ before $T_{c}$ in (d) and (e) (see also Refs. [53, 54, 55, 56]) is consistent with the theoretical prediction. In principle, quantitative agreement with specific experimental systems can be achieved by tuning the Josephson parameters ($J_{1}$, $J_{2}$, and $\alpha$) together with the disorder strength or carrier density.
We emphasize that the renormalized Josephson couplings in Eqs. (13) and (14) are derived in a mathematically rigorous manner, without reliance on model-specific parameters or physical assumptions, and are therefore generic. Then, the key point is whether Josephson phase fluctuations can be effectively excited. Typically, the Josephson coupling is weak, such that the gap of the Josephson phase mode $\omega_{L}(q=0)$ in Eq. (17)] lies below the SC gap, rendering it a thermally active low-energy excitation. Notably, an equally important factor is the phase stiffness $f_{s}$, which determines the available phase space for bosonic phase fluctuations entering the fluctuation integrals [Eq. (16)]. In bulk systems, the phase stiffness is generally large, strongly suppressing phase fluctuations. By contrast, in low-dimensional systems the phase stiffness is significantly reduced, owing both to a smaller Fermi velocity and to the intrinsically enhanced disorder associated with reduced dimensionality [57, 58]. Increasing disorder and reducing the Fermi velocity, as shown in Figs. 3(a) and (b), respectively, lead to a suppression of $T_{s}$, $T_{c}$, $T_{\eta}$, while markedly increasing the separation among these characteristic temperatures. As a result, the hierarchy of SC decoherence scales becomes increasingly pronounced. Among these effects, the suppression of gap closing temperature $T_{s}$ originates from the enhanced NG phase fluctuations (as discussed in Refs. [47, 48, 49]), whereas all the remaining Josephson features are primarily driven by the enhanced Josephson phase fluctuations (i.e., decoherence).
Discussion.—Beyond the conventional single-energy-scale paradigm of the Josephson physics, the present study demonstrates that phase decoherence constitutes an independent and more fragile energy scale in low-dimensional Josephson systems, and hence, Josephson physics itself in this case splits into multiple decoherence regimes, including diode nonreciprocity as the most fragile scale. We therefore predict that the forward and backward critical currents in low-dimensional Josephson-diode systems, $|I_{c}^{+}(T)|$ and $|I_{c}^{-}(T)|$, to merge as temperature increases, before vanishing altogether at $T_{c}$, as shown in Fig. 3(c). A similar tendency appears to be present in a wide range of recent experiments (e.g., Refs. [24, 23, 53, 54, 55, 56]), where the data [see Fig. 3(d) and (e)] exhibit a temperature evolution qualitatively consistent with theoretical prediction.
It should be emphasized that the vanishing of the Josephson critical current and the disappearance of the diode effect correspond to exponential thermal crossovers, rather than conventional thermodynamic phase transitions of first or second order. Both the diode efficiency $\eta(T)$ and the Josephson critical current $I_{c}(T)$ shown in Fig. 2(b) exhibit behaviors with continuous temperature derivatives in the vicinity of their respective disappearance temperatures, in contrast to the mean-field scenario shown in Fig. 2(a), where discontinuities in temperature derivatives are expected. Consequently, these crossovers do not exhibit thermodynamic signatures such as latent heat or a jump in the specific heat, in contrast to the SC gap-closing transition, which is a genuine second-order phase transition [33, 34]. Such critical behaviors of the Josephson properties thus constitute a key distinguishing feature between mean-field and fluctuation-based descriptions of Josephson systems. The proposed smooth-SC-decoherence mechanism follows analytically from the structure of both quantum and thermal phase fluctuations in low dimensions and does not rely on microscopic pairing details, symmetry-specific mechanisms, or fine tuning. It is fundamentally distinct from the BKT transition, the well-known mechanism for destroying SC phase coherence in two dimensions, which acts on the gapless global phase through vortex unbinding. By contrast, the Josephson phase mode is gapped due to interlayer phase locking and therefore does not participate in BKT physics.
The developed theory provides a tractable description of SC decoherence processes. Owing to the structural similarity [37, 49], this framework can be potentially extended to layered superconductors with multilayer structures and Josephson coupling $J_{c}\cos(\theta_{n+1}-\theta_{n})$ between different layers (where $\theta_{n}$ is the SC phase in the $n$-th layer), such as cuprates [59, 60, 61, 62, 63, 64, 65, 66] and recently discovered nickelates [67, 68, 69, 70, 71, 72], particularly for their $c$-axis transport behavior. In this context, one can anticipate the disappearance of $c$-axis superconductivity (zero-resistance) while the SC gap persists above a characteristic critical temperature. Moreover, owing to the close analogy in the underlying derivation of the excitations, similar considerations also apply to SC transmon qubits [73, 74, 75]. Our results indicate that higher-order Josephson harmonics are increasingly fragile against decoherence. This effect is expected to be relevant for the stability/coherence of Josephson-based SC qubits, where such higher harmonics play an essential role in determining the qubit spectrum and anharmonicity.
Acknowledgments.—This work is supported by the US Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of Computational Materials Sciences Program. F.Y. and L.Q.C. also appreciate the generous support from the Donald W. Hamer Foundation through a Hamer Professorship at Penn State.